OpenStudy (anonymous):

y=9-x^2,y=0, find volume for y=9

OpenStudy (anonymous):

y=9 is axis to rotate over

OpenStudy (anonymous):

The radius is given by 9-(9-x^2)=x^2 And differential of volume dV=PI*R^2*dx=PI*x^4*dx Limts of the integral are 9-x¨2=0->> x=-3 and x=3 \[V=π \int\limits_{-3}^{+3}x^{4} dx=π[3^{5} −(−3)^{5} ]/5=2π3^{5}/5=305.36UoV \]